princeevee:

Given that sinθ=35 , what is cosθ?

Hero:

@princeevee do you know the relationship between sin theta and cosine theta?

princeevee:

not really, i kinda forget stuff, can you refresh me on it?

Hero:

Suppose theta represents an angle of a right triangle: 

Hero:

Then \(\sin(\theta) = \dfrac{a}{c}\) and \(\cos(\theta) = \dfrac{b}{c}\) \(a = c\sin(\theta)\) and \(b = c\cos(\theta)\) And then you use pythagorean theorem at this point: \(a^2 + b^2 = c^2\) Which means: \((c\sin(\theta))^2 + (c\cos(\theta))^2= c^2\) which simplifies to: \(c^2\sin^2(\theta) + c^2(\cos^2(\theta) = c^2\) Factor out the \(c^2\) on the LHS to get: \(c^2(\sin^2(\theta) + \cos^2(\theta)) = c^2\) Divide both sides by \(c^2\) to get \(\sin^2(\theta) + \cos^2(\theta) = \dfrac{c^2}{c^2}\) or simply: \(\sin^2(\theta) + \cos^2(\theta) = 1\)

Hero:

We know \(\sin(\theta) = 35\) so we insert that in to the formula to get: \(35^2 + \cos^2(\theta) = 1\) The continue solving for \(\cos(\theta)\) BTW, \(\cos^2(\theta)\) and \((\cos(\theta))^2\) mean the same thing\

Hero:

Any questions @princeevee ?

princeevee:

i'm still a bit confused..

Hero:

About what?

princeevee:

the final answer is a fraction, right?

princeevee:

so do i need to have a number to place the cosine theta to complete the equation?

Hero:

We have to isolate \(\cos(\theta)\) in the equation \(35^2+ \cos^2(\theta) = 1\)

Hero:

What must we do first in order to accomplish this?

princeevee:

i...dont remember..

Hero:

Subtract 35^2 from both sides. What do you get?

princeevee:

so 35^2 - 35^2 and cos(theta) - 35^2?

Hero:

It's \(\sin(\theta) = 35\) not \(\theta= 35\)

charge69:

nvm, I read that wrong sorry to confuse you.

charge69:

Yea I noticed ty srry to interrupt

princeevee:

is my way of subtracting right?

Hero:

Do you know how to subtract a number from both sides?

Hero:



princeevee:

that's -1224, right?

Hero:

Correct and remember what I said earlier. You can't take the square root of a negative number.

Hero:

So what does this mean in terms of cosine theta?

princeevee:

the theta is positive?

Hero:

if \(\cos^2(\theta) = -1224\) what is the next step?

princeevee:

i dont know then..

Hero:

Have you even taken the square root of both sides?

Hero:

of an equation?

princeevee:

i did, it's just this is all confusing..

princeevee:

@Hero ?

Hero:

Sorry to hear that.

Hero:

The point is, you can't take the square root of a negative number and expect a real number result.

princeevee:

so what do we do to make it a positive?

Hero:

If a number is negative, then it is negative. You can't make a negative number positive.

princeevee:

so something went wrong with the subtraction?

Hero:

if you look, \(\cos(\theta) = b/c\) what is \(b/c\) ?

princeevee:

opposite / hypotenuse?

princeevee:

wait, i wrote something wrong!

princeevee:

it's 3/5, not 35

Hero:

smh

princeevee:

i know, sorry..

Hero:

This is why I don't help users with questions. I end up wasting my own time.

princeevee:

....

Hero:

You might as well just post a new question.

princeevee:

so i just post the same one with the correct changes?

Hero:

exactly